I HAVE THE FOLLOWING QUESTIONS IN MOTION IN A STRAIGHT LINE. I WOULD LIKE TO HAVE THE ANSWER FOR VERIFICATION.


 


QUESTION:


A baloon is ascending at the rate of 14 m/s at a height of 98 m above the ground, when a food packet is dropped from the baloon. After how much time and with what velocity does it reach the ground. Take g = 9.8 m / s2


 

Grade:11

5 Answers

sanatan sharma
11 years ago

applying the equation -   s=ut+(1/2)at2

here s = 98m

u = -14m/s

solving the equation we get 

98 = -14t + (1/2)9.8t2

4.9t2 - 14t - 98 = 0

solve the equation and the positive time will be the answer


Gouthuam Lakshminarayanan
11 years ago

YES. I liked the answer but my Physics teacher did not give me marks since I solved it using Quadriatic equations.

 

neeraj agarwal
11 years ago

ha gautam ji you should first know that if we drop packet from the ballon the packet will get the velocity of ballon at that instant and now                                                                                                                                                                  s=ut=1/2at2                                                                                                                                                                     -98=14t-1/2*9.8*t2                                                                                                                                                           get t and now use                                                                                                                                                             v=u+at          get velocitySmile

haritha m
9 years ago
-98 =14 * t – ( (9.8 * t2 )/ 2 )
on solving get t=3.3
 v same as the ballon
 
 
 
 
 
 
 
 
 
KAPIL MANDAL
9 years ago
Thank you sir for this answer. Thank you very much.